OpenStudy (anonymous):

perform the indicated operation: x^3-8/x^2+2x+4 ÷ (x^2-4)

OpenStudy (aroub):

okay u just have to factor and then divide!

OpenStudy (anonymous):

still lost

OpenStudy (anonymous):

note that \[8 = 2^3\]

OpenStudy (anonymous):

\[4=2^2\]

OpenStudy (anonymous):

\[x^3-8 = x^3 -2^3\] \[x^2-4 = x^2 -2^2\]

OpenStudy (anonymous):

ok i got 2/x+1

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

ahh i dont know i think i factored wrong

OpenStudy (anonymous):

you need to know the formulas for \[a^2-b^2 \] and \[a^3-b^3\] google them if you don't know the formulas already.

OpenStudy (anonymous):

\[\frac{x^3-8}{x^2+2x+4} ÷ (x^2-4)\] \[=\frac{(x-2)(x^2+2x+4)}{(x^2+2x+4)} \div (x-2)(x+2)\] \[=\frac{(x-2)(x^2+2x+4)}{(x^2+2x+4)} \times \frac{1}{(x-2)(x+2)}\] \[=\frac{1}{(x+2)}\]

OpenStudy (anonymous):

hakkunamatata, we generally try to get the students to work on their problems themselves. We don't do their problems for them.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

right

OpenStudy (anonymous):

whywhywhy!!!!